# Magnetic monopoles and high frequency gravitational waves from quasi-stable strings

### 03 Mar 2026 v1

@@ 1 Introduction
Dirac, nearly a century ago, postulated the existence of magnetic monopoles in order to explain electric charge quantization [1]. Spontaneously broken gauge theories with quantized electric charge predict the existence of topologically stable magnetic monopoles [2, 3]. In the framework of grand unified theories such as $SU(5)$ [4] and SO(10) [5, 6], the minimal monopole is superheavy and carries a single unit ($2\pi/e$) of Dirac magnetic charge, as well as some color magnetic charge which is screened [7, 8, 9]. Unified theories based on a gauge symmetry such as $SU(4)_{c}\times SU(2)_{L}\times SU(2)_{R}$ [10] predict the existence of a magnetic monopole that carries two units ($4\pi/e$) of Dirac charge as well as color magnetic charge [11, 12]. The mass of this monopole can vary over a wide range, as discussed recently [13, 14]. For a discussion on how topologically stable monopoles can arise from the decay of a metastable string network, see Ref. [13]. A discussion of how superheavy monopole may survive primordial inflation can be found in Ref. [15]. For recent studies on metastable and superconducting strings in $SO(10)$ and $E_{6}$, see Refs. [16, 17, 18, 19]. High frequency gravitational waves radiated from composite topological structures such as ‘walls bounded by strings’ [20] along with an observable flux of GUT monopoles, have been discussed in Ref. [21]. For other related studies see Refs. [22, 23, 24, 25, 26, 27, 28, 29].
@@ 3 Quasi-stable string and high frequency gravitational waves
Figure 5: Gravitational wave background from the quasi-stable string network for $G\mu=10^{-8}$, $10^{-7}$, and $10^{-6}$ with a varying monopole horizon re-entry time $t_{M}=10^{-16}$, $10^{-10}$, $10^{-4}$ and $10^{2}$ sec. We also display the power-law integrated sensitivity curves [62, 63] for planned experiments near Hz to kHz frequency region, namely, HLVK [64], ET [65], CE [66], DECIGO [67], BBO [68, 69], LISA and SKA. The gray shaded region shows the bound for a scale-invariant gravitational wave spectrum for $f_{\rm high}/f_{\rm low}=10^{7}$ arising from the bound on $\Delta N_{\rm eff}$ in the BBN and CMB data.
@@
It is worth mentioning at this point that the quasi-stable strings can explain the evidence of a gravitational wave background in the recent NANOGrav 15 year data for $G\mu\in[2\times 10^{-8},7\times 10^{-6}]$ and $t_{M}\in[2\times 10^{1},1\times 10^{5}]$ sec (95% Bayesian credible intervals) [73, 74]. The gravitational waves for $G\mu=[10^{-8},10^{-7}]$ can be detected in all the proposed experiments depending on the horizon re-entry time of the monopoles. The constraint from the third observing run data from LIGO and VIRGO experiments for $G\mu>10^{-7}$ can be alleviated if the strings experience a certain number of $e$-foldings during inflation for high $t_{M}$ values. For example, $G\mu=10^{-6}$ and $t_{M}=10^{2.5}$ sec can explain the NANOGrav data and satisfy the LV3 bound for $t_{s}=10^{-10}$ sec, as shown in Ref. [74].

### 05 Mar 2026 v2

@@ 1 Introduction
Dirac, nearly a century ago, postulated the existence of magnetic monopoles in order to explain electric charge quantization [1]. Spontaneously broken gauge theories with quantized electric charge predict the existence of topologically stable magnetic monopoles [2, 3]. In the framework of grand unified theories such as $SU(5)$ [4] and SO(10) [5, 6], the minimal monopole is superheavy and carries a single unit ($2\pi/e$) of Dirac magnetic charge, as well as some color magnetic charge which is screened [7, 8, 9]. Unified theories based on a gauge symmetry such as $SU(4)_{c}\times SU(2)_{L}\times SU(2)_{R}$ [10] predict the existence of a magnetic monopole that carries two units ($4\pi/e$) of Dirac charge as well as color magnetic charge [11, 12]. The mass of this monopole can vary over a wide range, as discussed recently [13, 14]. For a discussion on how topologically stable monopoles can arise from the decay of a metastable string network, see Ref. [13]. A discussion of how superheavy monopole may survive primordial inflation can be found in Ref. [15]. For recent studies on metastable and superconducting strings in $SO(10)$ and $E_{6}$, see Refs. [16, 17, 18, 19]. High frequency gravitational waves radiated from composite topological structures such as ‘walls bounded by strings’ [20] along with an observable flux of GUT monopoles, have been discussed in Ref. [21]. For other related studies see Refs. [22, 23, 24, 25, 26, 27, 28, 29, 30]. For a discussion of the primordial monopole problem, see Refs. [31, 32].
@@ 3 Quasi-stable string and high frequency gravitational waves
Figure 5: Gravitational wave background from the quasi-stable string network for $G\mu=10^{-8}$, $10^{-7}$, and $10^{-6}$ with a varying monopole horizon re-entry time $t_{M}=10^{-16}$, $10^{-10}$, $10^{-4}$ and $10^{2}$ sec. We also display the power-law integrated sensitivity curves [65, 66] for planned experiments near Hz to kHz frequency region, namely, HLVK [67], ET [68], CE [69], DECIGO [70], BBO [71, 72], LISA [73, 74] and SKA [75]. The gray shaded region shows the bound for a scale-invariant gravitational wave spectrum for $f_{\rm high}/f_{\rm low}=10^{7}$ arising from the bound on $\Delta N_{\rm eff}$ in the BBN and CMB data. The red violin plots represent the Bayesian free-spectral periodogram from the NANOGrav 15 year data [76, 77].
@@
It is worth mentioning at this point that the quasi-stable strings can explain the evidence of a gravitational wave background in the recent NANOGrav 15 year data [76, 77] for $G\mu\in[2\times 10^{-8},7\times 10^{-6}]$ and $t_{M}\in[2\times 10^{1},1\times 10^{5}]$ sec (95% Bayesian credible intervals) [81, 82]. The gravitational waves for $G\mu=[10^{-8},10^{-7}]$ can be detected in all the proposed experiments depending on the horizon re-entry time of the monopoles. The constraint from the third observing run data from LIGO and VIRGO experiments for $G\mu>10^{-7}$ can be alleviated if the strings experience a certain number of $e$-foldings during inflation for high $t_{M}$ values. For example, $G\mu=10^{-6}$ and $t_{M}=10^{2.5}$ sec can explain the NANOGrav data and satisfy the LV3 bound for $t_{s}=10^{-10}$ sec, as shown in Ref. [82].